OpenStudy (anonymous):

(1)/(5)=()/(15) find the number with denominator indicated multiply by 1

OpenStudy (anonymous):

1/5 = 3/15

OpenStudy (anonymous):

3

OpenStudy (anonymous):

yes, correc

OpenStudy (anonymous):

yes thats write

OpenStudy (anonymous):

yhnaks thanks

OpenStudy (anonymous):

welcome...fan me please

OpenStudy (anonymous):

(5)/(7)=()/(49) find the number with denominator indicated multiply by 1

OpenStudy (anonymous):

35

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

do u understand...equivalent fractions

OpenStudy (anonymous):

using multiplying by 1 to find and expression equivalent (17) (14) with a denominator of 14d

OpenStudy (anonymous):

what would be my equivalte exression be

OpenStudy (anonymous):

17*14=17/(1/14)

OpenStudy (anonymous):

divide fractions-multiply by the reciprocal (multiplicative inverse)-flip numerator and denominator

OpenStudy (anonymous):

so the expression will what 17x14=17 /(1/14)

OpenStudy (anonymous):

yes : )

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

17/(1/14)

OpenStudy (anonymous):

nope it d17 d114

OpenStudy (anonymous):

d17 d14

OpenStudy (anonymous):

using multiplying by 1 to find and expression equivalent (11) (15) with a denominator of 15x

OpenStudy (anonymous):

what would be my equivalte exression be